# Models Got Talent: Identifying High Performing Wearable Human Activity Recognition Models Without Training

### v1

@@ Abstract
Discovering high performing model architectures for wearables-based Human Activity Recognition (HAR) applications is challenging. The astonishing diversity and variability due to differing sensor locations, recording apparatus, activities etc., can cause established baseline architectures to perform worse on datasets/tasks they were not originally trained on. A promising alternative to the computationally expensive Neural Architecture Search (NAS) involves the development of Zero Cost Proxies (ZCPs), which correlate well to trained performance, but can be computed through a single forward/backward pass on a randomly sampled batch of data. In this paper, we investigate the effectiveness of ZCPs for HAR on six benchmark datasets, and demonstrate that they discover network architectures that obtain within 5% of performance attained by full scale training involving 1500 randomly sampled architectures. This results in substantial computational savings as high performing architectures can be discovered with minimal training. Our experiments not only introduce ZCPs to sensor-based HAR, but also demonstrate that they are robust to data noise, further showcasing their suitability for practical scenarios.
@@ V-B Talent Rates of ZCPs
We perform full training with the top-10% of models predicted by ZCPs and denote the difference in performance to the top model as $\Delta_{10\%}$. We observe that $\Delta_{10\%}<\Delta_{1}$ overall and most ZCP produce $\Delta_{10\%}<1\%$. Similar to Fig.˜2, the random search can also be effective, but also suffers from susbtantial variance.

### v2

@@ Abstract
Discovering high performing model architectures for wearables-based Human Activity Recognition (HAR) applications is challenging. The astonishing diversity and variability due to differing sensor locations, recording apparatus, activities etc., can cause established baseline architectures to perform worse on datasets/tasks they were not originally trained on. A promising alternative to the computationally expensive Neural Architecture Search (NAS) involves the development of Zero Cost Proxies (ZCPs), which correlate well with trained performance, but can be computed through a single forward/backward pass on a randomly sampled batch of data. In this paper, we investigate the effectiveness of ZCPs for HAR on six benchmark datasets, and demonstrate that they discover network architectures that obtain within 5% of performance attained by full-scale training involving 1500 randomly sampled architectures. This results in substantial computational savings as high-performing architectures can be discovered with minimal training. Our experiments not only introduce ZCPs to sensor-based HAR, but also demonstrate that they are robust to data noise, further showcasing their suitability for practical scenarios.
@@ V-B Talent Rates of ZCPs
We perform full training with the top-10% of models predicted by ZCPs and denote the difference in performance to the top model as $\Delta_{10\%}$. We observe that $\Delta_{10\%}<\Delta_{1}$ overall and most ZCP produce $\Delta_{10\%}<1\%$. Similar to Fig.˜2, the random search can also be effective, but also suffers from substantial variance.